A person has thirty coins, comprised of nickles and dimes. What is the maximum number of nickels if the total value is at most $2.60?
Answer by josgarithmetic(39633)

d dimes and n nickels.
d+n=30 and 0.1d+0.05n=2.60.
You are not looking for a maximum quantity, but are looking for an exact quantity, unless it is not a whole number.

Find what happens.
d+n=30 and 2d+n=52

2d+n-(d+n)=52-30
d=22
meaning that n=8.